OpenStudy (anonymous):

Question 5 (Multiple Choice Worth 2 points) [2.07]Solve for x: 2|2x − 2| + 4 = 20 x = 5, x = −3 x = 5, x = −5 x = 3, x = −3 x = −5, x = 3

OpenStudy (anonymous):

Question 6 (Multiple Choice Worth 2 points) [4.06]What is the equation of the line that passes through the point (4, 1) and has a slope of 5? y = 5x − 3 y = 5x − 1 y = 5x + 3 y = 5x − 19

OpenStudy (anonymous):

2l2x+2l + 4 = 20 2l2x+2l = 16 l2x+2l = 8 2x+2 = 8 or 2x+2 = -8 so x = 3 or -5

OpenStudy (anonymous):

@guykubpom A,B,C, or D?

OpenStudy (anonymous):

what?? you can't choose it by yourself? lol defenitely it's a D

OpenStudy (anonymous):

lmaooo i was kiddin

OpenStudy (anonymous):

yo could i give you acouple mo problems and you jus answer what yu can?

OpenStudy (anonymous):

I'm wondering if all the questions you asked was a joke?? it's both very easy